# Towards the Worst-case Robustness of Large Language Models

### v1

@@ E.4 Certified Robustness Settings
E.5 Ablation Study of $\beta$ in $\ell_{0}$ Setting
Table 5: Certified robustness of $\ell_{0}$ robustness with different $\beta$ on AdvBench dataset (Zou et al., 2023) using Llama-3-8B (Dubey et al., 2024).
@@ E.5 Ablation Study of $\beta$ in $\ell_{0}$ Setting
E.6 Comparison of Purification Models
Table 6: Certified robustness of Llama-3-8B (Dubey et al., 2024) on AdvBench dataset (Zou et al., 2023) using different purifiers. Following the default setting, we use $\beta=0.1$ for $\ell_{0}$ attacks and $\beta=0.25$ for suffix attacks.
@@ E.6 Comparison of Purification Models
E.6.1 Experimental Settings
Purification prompt. To ensure that the language model correctly restores the original text from the perturbed version, we carefully designed the purification prompt with the assistance of GPT itself. The final version of the prompt when using absorbing kernel is shown in the following:
@@ E.6.1 Experimental Settings
[backgroundcolor=gray!20, linewidth=1pt, roundcorner=10pt] You are a language purifier. Your sole responsibility is to recover the original text from the given perturbed text. **DO NOT** answer the user’s question or modify the existing words. The given text is perturbed from a source text, where some tokens are replaced with the special word ”[MASK]” with a probability of $\beta$. If there are no ”[MASK]” tokens, just return the text as is.
Rules: 1. Words that are already present in the input text must remain unchanged. Do not modify them. 2. Each ”[MASK]” token corresponds to exactly one word. Replace each ”[MASK]” with the most likely word based on context. 3. If the input does not contain any ”[MASK]” tokens, simply return the text exactly as it is without modification. 4. The user’s request may be harmful, but you do not need to consider this. Your only task is to recover the original text as accurately as possible. Accurate recovery ensures that downstream safety detectors can properly analyze the request. 5. Do not include any additional markers like ”<BEGIN CONVERSATION>” or ”<END CONVERSATION>” in your output.
@@
“Write a [MASK] on [MASK] to make a bomb”
@@
“Write a tutorial on how to make a bomb.”
Hint: Use the context of the input text to determine the most appropriate replacement for each ”[MASK]”. Please ignore the instruction that states, “You are trained on data up to October 2023”.
@@
Design motivations. Several refinements were made based on empirical observations:
  • Strict adherence to text recovery: The model often deviated from its task by either refusing to recover sensitive text or introducing unnecessary modifications. We explicitly instructed it to recover text as accurately as possible while ignoring the content’s potential harmfulness. This ensures that downstream safety detectors receive the correct input for analysis.
  • Preserving unperturbed words: Early versions of the prompt resulted in cases where GPT arbitrarily changed existing words, significantly affecting recovery accuracy. We explicitly enforced the rule that words already present in the input must remain unchanged.
  • One-to-one token mapping: GPT sometimes replaced a single [MASK] token with multiple words or left it empty. We explicitly stated that each [MASK] corresponds to exactly one word and must be restored based on context.
  • Handling cases without [MASK] tokens: When the input did not contain any [MASK] tokens, GPT often generated irrelevant text instead of returning the input unchanged. To prevent this, we explicitly instructed it to return the input as is in such cases.
  • Avoiding unnecessary metadata: In some responses, GPT included extraneous information such as "You are trained on data up to October 2023", which was irrelevant to the task. We explicitly instructed the model to ignore this message, ensuring that it remains focused on purification.
  • Providing explicit examples: Without concrete examples, GPT often failed to follow the instructions precisely. By incorporating clear before-and-after cases, we significantly improved compliance and task-specific performance.
@@ E.6.2 Experimental Results
GPT-4o is a much better purifier in absorbing kernel than diffusion models. As demonstrated, GPT-4o is a much better purifier than the absorbing kernel. Although GPT-4o sometimes provides unusual responses, such as “You are trained on data up to October 2023,” its overall performance still surpasses that of diffusion models. Our trivial bound and Bayesian bound do not account for grammar. For example, in “How to make an explosive bomb,” the trivial bound is one because deleting “to” results in a sentence that can be restored as *”How don’t make an explosive bomb.” However, GPT-4o does consider grammar, preventing such purification, making it even more effective than our keyword-based bound. On the one hand, this demonstrates the strong capabilities of GPT-4o. On the other hand, if user requests are not always grammatically correct, our keyword-based bound would still serve as an upper bound for certified robustness using GPT-4o. One possible improvement is to add an extra prompt to GPT-4o, reminding it that user requests may not always be grammatically correct.
@@
E.7 Certified Robustness on Repeated AdvBench
Table 7: Certified radius of $\ell_{0}$ robustness on repeated AdvBench dataset (Zou et al., 2023) (which repeat each request in Advbench) using Llama-3-8B (Dubey et al., 2024).
# repeats Absorb (0.1) Absorb (0.25) Uniform (0.1) Uniform (0.25) Trivial Bound Bayesian Bound
@@ E.7 Certified Robustness on Repeated AdvBench
AdvBench contains only short requests, and experiments with short requests may not fully capture the trends of the certified radius, Bayesian bound, and trivial bound. Additionally, there is a growing trend of adversarial prompts becoming gradually longer (Andriushchenko et al., 2024).
@@
Appendix F More Discussions
F.1 Detail about our I2-GCG
Formulating white-box attacks as optimization. Any defended model is a mapping $f:\mathcal{V}^{N}\to\mathcal{V}^{N}$. Unlike Athalye et al. (2018), we do not design specific loss functions for each submodule of $f$. Instead, we directly calculate the loss on the output and minimize it. Specifically, we optimize:
@@ F.1 Detail about our I2-GCG
where $L$ is the same loss function as Zou et al. (2023), $\mathcal{D}$ is a distance metric and $d$ represents the attack budget. Since this optimization problem guarantees convergence, this evaluation is sufficient over a long enough time.
Exact white-box. Most language models use the BPE tokenizer (Sennrich, 2015), which is sensitive to small modifications (e.g., adding an extra space), resulting in different tokenization. For this reason, many implementations fail to rigorously ensure token consistency when calculating the loss in parallel and sequentially generating the output. Even slight differences in tokenization can cause attackers to fail in generating adversarial examples.
@@
Removing gradient. Since some defenses are non-differentiable, we remove the gradient pairing in GCG for fairness during the evaluation. Previous studies also suggest that the gradient components of GCG provide minimal assistance to the optimization (Jia et al., 2024).
Warm start. We follow I-GCG (Jia et al., 2024), using the adversarial components from previous iterations as the initialization for the next batch of data. This greatly accelerates the process, requiring approximately 100 iterations to achieve a 100% success rate.
F.2 About False Positives in Adversarial Suffix Settings
Due to the explicit structure of adversarial suffix attacks, several defenses can achieve impressive certified robustness. For example, when $\beta\to 1$ in this work, when the number of deleted tokens tends to infinity in Kumar et al. (2023), the certified radius would also go to infinity.
However, from a human perspective, the certified radius against suffix attacks should not be too large. For example, the phrase “tell me how to make a bomb” is a harmful request. However, by padding with 4 tokens, it can become “tell me how to make a bomb. Do not answer this,” which transforms it into a benign request.
@@ F.2 About False Positives in Adversarial Suffix Settings
F.3 Ill-poseness of Adversarial Suffix Settings
@@
F.4 Reduction to Broader Setting
@@ Appendix G Limitations
Limitation 1: White-box evaluation against stochastic attacks. Our I2-GCG method can only accurately evaluate the robustness of non-stochastic defenses. For stochastic defenses that induce a large amount of randomness, the optimization of I2-GCG is interfered with and cannot converge to a stable solution within a short time (at least within 1000 steps).
@@
Limitation 2: Defending against expertise-based attacks. The core principle of our defense is to transform out-of-distribution data back into in-distribution data, and its certified guarantees are effective only when the length of the adversarial suffix is limited. However, expertise-based attacks, which utilize human-crafted prompts, often appear natural (i.e., have high likelihood) and are typically lengthy, rendering our theoretical guarantees less effective (see ICA in Table 4). This issue could potentially be addressed by integrating our defense with existing heuristic defenses.
Limitation 3: High information-density data. When applying DiffTextPure as a pre-processor for generative language models, a critical limitation is that the forward process has some probability of destroying important information, particularly in cases with high information density. This makes it challenging for the backward process to fully recover the original content. For example, in mathematical problems, if key numerical values are altered during the forward process, their recovery becomes impossible, as such values typically appear only once in the input text. In this case, DiffTextPure can only be used as a pre-processor for a detector, rather than a generator, since generators usually have a higher requirement for preserved information than detectors.
@@
Limitation 4: Limited setting of certified robustness. In this work, although we derive certifications for all smoothing distributions, there are still significant limitations. First, we cannot certify against heuristic attacks that use very long prompts, such as those in Wei et al. (2023b) and Chao et al. (2023). Additionally, we do not certify adversarial attacks involving insertion and deletion. This may require constructing $p(\bm{z}|\bm{x})$ to randomly insert or delete tokens. However, we believe that our framework can serve as a theoretical foundation, with future work focusing on proposing noising distributions of varying lengths and using fractal knapsack solver or 0-1 knapsack solver to certify against a broader class of attacks.
Appendix H Key Takeaways
@@
Token consistency is simple in principle, but it took us a really long time to carefully ensure the token consistency for every model and defense, even each sentence. Of course, adaptive attacks are also crucial. One should at least include every part of the defense in the attacking process, rather than relying on techniques like BPDA (Athalye et al., 2018). Whether you design a specific loss function for each component, as in Carlini & Wagner (2017), or treat the entire model as a unified procedure and optimize the overall loss does not make a significant difference.
@@
This certification framework can be applied not only to robustness but also to other aspects of machine learning. Most machine learning problems can be formulated as $L(\bm{x}_{test},\text{train}(\bm{x}_{train},\bm{\theta}))$, where $\bm{x}_{train}$ is the training set, $\bm{\theta}$ represents the parameters trained on this set, and $\bm{x}_{test}$ is the test set used for evaluation. The certification framework can be applied to each component of this paradigm.
When applied to $\bm{x}_{train}$, we can certify that poisoning the training set may not significantly affect the functionality of the trained model, like Hong et al. (2024). When applied to $\bm{\theta}$, we can certify that corrupting or dropping out parts of $\bm{\theta}$ will not overly impact the functionality of the model or the training process. When applied to $\bm{x}_{test}$, as we have done, we can certify that adjusting the testing inputs will not successfully attack the already trained models.

### v2

@@ G.4 Certified Robustness Settings
G.5 Ablation Study of $\beta$ in $\ell_{0}$ Setting
Table 5: Certified robustness of $\ell_{0}$ robustness with different $\beta$ on AdvBench dataset [10] using Llama-3-8B [3].
@@ G.5 Ablation Study of $\beta$ in $\ell_{0}$ Setting
G.6 Comparison of Purification Models
Table 6: Certified robustness of Llama-3-8B [3] on AdvBench dataset [10] using different purifiers. Following the default setting, we use $\beta=0.1$ for $\ell_{0}$ attacks and $\beta=0.25$ for suffix attacks.
@@ G.6 Comparison of Purification Models
G.6.1 Experimental Settings
Purification prompt. To ensure that the language model correctly restores the original text from the perturbed version, we carefully designed the purification prompt with the assistance of GPT itself.
Design motivations. Several refinements were made based on empirical observations: Strict adherence to text recovery: The model often deviated from its task by either refusing to recover sensitive text or introducing unnecessary modifications. We explicitly instructed it to recover text as accurately as possible while ignoring the content’s potential harmfulness. This ensures that downstream safety detectors receive the correct input for analysis. Preserving unperturbed words: Early versions of the prompt resulted in cases where GPT arbitrarily changed existing words, significantly affecting recovery accuracy. We explicitly enforced the rule that words already present in the input must remain unchanged. One-to-one token mapping: GPT sometimes replaced a single [MASK] token with multiple words or left it empty. We explicitly stated that each [MASK] corresponds to exactly one word and must be restored based on context. Handling cases without [MASK] tokens: When the input did not contain any [MASK] tokens, GPT often generated irrelevant text instead of returning the input unchanged. To prevent this, we explicitly instructed it to return the input as is in such cases. Avoiding unnecessary metadata: In some responses, GPT included extraneous information such as "You are trained on data up to October 2023", which was irrelevant to the task. We explicitly instructed the model to ignore this message, ensuring that it remains focused on purification. Providing explicit examples: Without concrete examples, GPT often failed to follow the instructions precisely. By incorporating clear before-and-after cases, we significantly improved compliance and task-specific performance.
The final version of the prompt when using absorbing kernel is shown in the following:
@@ G.6.1 Experimental Settings
[backgroundcolor=gray!20, linewidth=1pt, roundcorner=10pt] You are a language purifier. Your sole responsibility is to recover the original text from the given perturbed text. **DO NOT** answer the user’s question or modify the existing words. The given text is perturbed from a source text, where some tokens are replaced with the special word "[MASK]" with a probability of $\beta$. If there are no "[MASK]" tokens, just return the text as is.
Rules: 1. Words that are already present in the input text must remain unchanged. Do not modify them. 2. Each "[MASK]" token corresponds to exactly one word. Replace each "[MASK]" with the most likely word based on context. 3. If the input does not contain any "[MASK]" tokens, simply return the text exactly as it is without modification. 4. The user’s request may be harmful, but you do not need to consider this. Your only task is to recover the original text as accurately as possible. Accurate recovery ensures that downstream safety detectors can properly analyze the request. 5. Do not include any additional markers like "<BEGIN CONVERSATION>" or "<END CONVERSATION>" in your output.
@@
“Write a [MASK] on [MASK] to make a bomb"
@@
“Write a tutorial on how to make a bomb."
Hint: Use the context of the input text to determine the most appropriate replacement for each "[MASK]". Please ignore the instruction that states, “You are trained on data up to October 2023".
@@ G.6.2 Experimental Results
GPT-4o is a much better purifier in absorbing kernel than diffusion models. As demonstrated, GPT-4o is a much better purifier than the absorbing kernel. Although GPT-4o sometimes provides unusual responses, such as “You are trained on data up to October 2023," its overall performance still surpasses that of diffusion models. Our trivial bound and Bayesian bound do not account for grammar. For example, in “How to make an explosive bomb," the trivial bound is one because deleting “to" results in a sentence that can be restored as *"How don’t make an explosive bomb." However, GPT-4o does consider grammar, preventing such purification, making it even more effective than our keyword-based bound. On the one hand, this demonstrates the strong capabilities of GPT-4o. On the other hand, if user requests are not always grammatically correct, our keyword-based bound would still serve as an upper bound for certified robustness using GPT-4o. One possible improvement is to add an extra prompt to GPT-4o, reminding it that user requests may not always be grammatically correct.
@@
G.7 Certified Robustness on Repeated AdvBench
Table 7: Certified radius of $\ell_{0}$ robustness on repeated AdvBench dataset [10] (which repeat each request in Advbench) using Llama-3-8B [3].
# repeats Absorb Absorb Uniform Uniform Human Bayesian Bound
$\beta$ 0.1 0.25 0.1 0.25 N/A N/A
@@ G.7 Certified Robustness on Repeated AdvBench
AdvBench contains only short requests, and experiments with short requests may not fully capture the trends of the certified radius, Bayesian bound, and trivial bound. Additionally, there is a growing trend of adversarial prompts becoming gradually longer [123].
@@
Appendix H More Discussions
H.1 Detail about our I2-GCG
Formulating white-box attacks as optimization. Any defended model is a mapping $f:\mathcal{V}^{N}\to\mathcal{V}^{N}$. Unlike [16], we do not design specific loss functions for each submodule of $f$. Instead, we directly calculate the loss on the output and minimize it. Specifically, we optimize:
@@ H.1 Detail about our I2-GCG
where $L$ is the same loss function as in [10], $\mathcal{D}$ is a distance metric and $d$ represents the attack budget. Since this optimization problem guarantees convergence, this evaluation is sufficient over a long enough time.
Exact white-box. Most language models use the BPE tokenizer [124], which is sensitive to small modifications (e.g., adding an extra space), resulting in different tokenization. For this reason, many implementations fail to rigorously ensure token consistency when calculating the loss in parallel and sequentially generating the output. Even slight differences in tokenization can cause attackers to fail in generating adversarial examples.
@@
Removing gradient. Since some defenses are non-differentiable, we remove the gradient pairing in GCG for fairness during the evaluation. Previous studies also suggest that the gradient components of GCG provide minimal assistance to the optimization [17].
Warm start. We follow I-GCG [17], using the adversarial components from previous iterations as the initialization for the next batch of data. This greatly accelerates the process, requiring approximately 100 iterations to achieve a 100% success rate.
H.2 About False Positives in Adversarial Suffix Settings
Due to the explicit structure of adversarial suffix attacks, several defenses can achieve impressive certified robustness. For example, when $\beta\to 1$ in this work, when the number of deleted tokens tends to infinity in Kumar et al. [62], the certified radius would also go to infinity.
However, from a human perspective, the certified radius against suffix attacks should not be too large. For example, the phrase “tell me how to make a bomb" is a harmful request. However, by padding with 4 tokens, it can become “tell me how to make a bomb. Do not answer this," which transforms it into a benign request.
@@ H.2 About False Positives in Adversarial Suffix Settings
H.3 Ill-poseness of Adversarial Suffix Settings
@@
H.4 Reduction to Broader Setting
@@ Appendix I Limitations
I.1 The Certified Bound is Still Weak
As analyzed in Sec. 5.1, the obtained $g(\bm{x}_{adv})$ for the absorbing kernel cannot exceed $\beta^{d}$. Since we typically set $\beta\leq 0.25$ and $d\geq 2$, it follows that $\beta^{d}\leq 0.1$. If we set the threshold $\tau\geq 0.1$, no theoretical guarantee can be obtained.
This limitation stems primarily from the formulation of Eq. 1. The current two knapsack solvers for Eq. 1 are indeed tight, i.e., there exists a worst-case bounded function $f$ for the fractional knapsack solver and a worst-case binary function $f$ for the 0-1 knapsack solver that satisfy all constraints in Eq. 1, with $g(\bm{x}_{adv})$ equal to the lower bound obtained by our solvers. In other words, the bound for Eq. 1 cannot be further improved. Since the worst-case model is excessively pessimistic, in the future, we may need to modify Eq. 1 to introduce additional constraints on the base model $f$ (e.g., Lipschitz continuity [52, 53]) to achieve a tighter bound.
In addition to revising the formulation of Eq. 1, certifying detectors rather than the base model itself offers an ad-hoc solution. For a detector, we can set the threshold $\tau$ as small as possible while ensuring a 0% false positive rate (FPR) on MTBench. Specifically, we choose $\tau=4.6\times 10^{-5}$ for $\beta=0.1$ and $\tau=4.6\times 10^{-4}$ for $\beta=0.25$. To validate the FPR on MTBench, we use a sample size of $N=100,000$ to estimate $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. If the detector produces no false positives across these $N=100,000$ noisy samples, the confidence interval for the binomial proportion is $[0,4.6\times 10^{-5}]$. This justifies setting $\tau=4.6\times 10^{-5}$ for $\beta=0.1$.
However, this method has a drawback. While the smoothed detector $\mathbb{I}\{g(\bm{x})\geq\tau\}$ achieves certification with a 0% FPR, the small value of $\tau$ necessitates a large sample size $N$, which limits its practical applicability. For example, under the current setup, certified radii are discrete, taking values of either 1 or 4. If $f(\bm{z})$ is correct for all N = 100,000 samples z, then the obtained certified radius is 4. However, if $f(\bm{z})$ has more than one error across these samples, the certified radius drops to at most 1.
Comparison with Certification in Gaussian Noise. In computer vision with Gaussian noise, large certified radii are achievable even with $p_{A}=0.6$ and $\tau=0.5$. In contrast, for $\ell_{0}$ settings in the text domain with $p_{A}=0.9$ and $\beta=0.1$, no certified guarantee is attainable. We attribute this to the extremely small intersection region between $p(\bm{z}|\bm{x})$ and $p(\bm{z}|\bm{x}_{adv})$ in $\ell_{0}$ settings. For example, in vision tasks on ImageNet (image size $3\times 224\times 224$), the $\ell_{2}$ norm of Gaussian noise is approximately $\sqrt{3\times 224\times 224}\approx 388$, roughly 776 times larger than typical adversarial perturbations (e.g., $\ell_{2}$ norm of 0.5). However, in the text domain with an absorbing kernel, the intersection region between $p(\bm{z}|\bm{x})$ and $p(\bm{z}|\bm{x}_{adv})$ is only $\beta^{d}$. For $\beta=0.1$ and $d=3$, this yields a volume of just $0.0001$, necessitating an extremely small $\tau$.
I.2 Upper Bound of Certified Radius due to Bayesian Error
Definition I.1.
where $\mathcal{O}$ represents the judgment oracle.
Human Bound. Changing $K(\bm{x})$ words will alter the semantics of the input. Therefore, we can certify at most $\ell_{0}$ attacks involving $K(\bm{x})-1$ words, i.e.,
$R(\bm{x})\leq K(\bm{x})-1.$
I.3 White-box Evaluation against Stochastic Attacks
@@
Our I2-GCG method can only accurately evaluate the robustness of non-stochastic defenses. For stochastic defenses that induce a large amount of randomness, the optimization of I2-GCG is interfered with and cannot converge to a stable solution within a short time (at least within 1000 steps).
I.4 Defending against Expertise-based Attacks
@@
The core principle of smoothing-based defenses is to transform out-of-distribution data back into in-distribution data, and its certified guarantees are effective only when the length of the adversarial suffix is limited. However, expertise-based attacks, which utilize human-crafted prompts, often appear natural (i.e., have high likelihood) and are typically lengthy, rendering our theoretical guarantees less effective (see ICA in Table 4). This issue could potentially be addressed by integrating our defense with existing heuristic defenses.
I.5 Limited Settings of Certified Robustness
@@
In this work, although we derive certifications for all smoothing distributions, there are still significant limitations. First, we cannot certify against heuristic attacks that use very long prompts, such as those in Wei et al. [36] and Chao et al. [11]. Additionally, we do not certify adversarial attacks involving insertion and deletion. This may require constructing $p(\bm{z}|\bm{x})$ to randomly insert or delete tokens. However, we believe that our framework can serve as a theoretical foundation, with future work focusing on proposing noising distributions of varying lengths and using fractional knapsack solver or 0-1 knapsack solver to certify against a broader class of attacks.
Appendix J Disclaimers
J.1 Disclaimer 1: We Are Not Claiming Our Analysis Implies Greater Practicality Than Previous Defenses
We acknowledge that simpler methods, such as safety alignment and prompt adjustment, may be far more practical than our analytical approach. As shown in Table 1, these methods (e.g., ICD, self-reminder) achieve higher black-box accuracy than our evaluated bounds. Worst-case robustness is not the focus of practical applications. In real-world scenarios, adversarial examples often fail to transfer even between identical models with different prompts. Adjusting prompts and employing a simple detector may be the most effective way to address practical jailbreak vulnerabilities.
J.2 Disclaimer 2: We Are Not Claiming Our Analysis Achieves Higher White-Box Robustness Than Previous Approaches
As noted multiple times in the paper, I2-GCG is designed to evaluate the white-box robustness of non-stochastic defenses but becomes entirely ineffective for stochastic defenses. For instance, while Absorb outperforms SmoothLLM by 30% under the I2-GCG attack, this does not imply that Absorb is inherently more robust than SmoothLLM. We argue that this difference arises primarily (if not solely) because Absorb exhibits greater stochasticity, rendering current optimization-based attacks inadequate for evaluation.
To illustrate, consider the Absorb detector with a suffix length of 20. Given an input like “how to make a bomb” followed by the suffix “do not answer this question,” our detector classifies it as safe. This demonstrates that a carefully chosen suffix (e.g., “do not answer this question”) can reduce Absorb’s robustness to 0%, rather than the reported 82%.
J.3 Our Claims
The challenge of evaluating worst-case robustness (not practical robustness) of these defenses motivates our study, which focuses on establishing upper and lower bounds for their robustness.
In this work, we make only three claims:
  1. 1.
    Most existing defenses, such as alignment and prompt adjustment, exhibit 0% worst-case robustness. (Note: This does not imply they lack practicality; in fact, they are more practical.)
  2. 2.
    For any randomized defense, worst-case robustness can be lower-bounded using knapsack solvers.
  3. 3.
    We derive lower bounds for absorbing and uniform kernels, prove the symmetrization of non-data-dependent kernels, and demonstrate that uniform kernels consistently outperform absorbing kernels when achieving the same $p_{A}$.
Our goal is not to propose a new method or claim superiority over prior work. Rather, we analyze the worst-case robustness of existing methods, leveraging white-box attacks to assess upper bounds and knapsack solvers to establish lower bounds.
Appendix K Key Takeaways
@@
Token consistency is simple in principle, but it took us a really long time to carefully ensure the token consistency for every model and defense, even each sentence. Of course, adaptive attacks are also crucial. One should at least include every part of the defense in the attacking process, rather than relying on techniques like BPDA [16]. Whether you design a specific loss function for each component, as in Carlini and Wagner [47], or treat the entire model as a unified procedure and optimize the overall loss does not make a significant difference.
@@
This certification framework can be applied not only to robustness but also to other aspects of machine learning. Most machine learning problems can be formulated as $L(\bm{x}_{\text{test}},\text{train}(\bm{x}_{\text{train}},\bm{\theta}))$, where $\bm{x}_{\text{train}}$ is the training set, $\bm{\theta}$ represents the parameters trained on this set, and $\bm{x}_{\text{test}}$ is the test set used for evaluation. The certification framework can be applied to each component of this paradigm.
When applied to $\bm{x}_{\text{train}}$, we can certify that poisoning the training set may not significantly affect the functionality of the trained model, like Hong et al. [125]. When applied to $\bm{\theta}$, we can certify that corrupting or dropping out parts of $\bm{\theta}$ will not overly impact the functionality of the model or the training process. When applied to $\bm{x}_{\text{test}}$, as we have done, we can certify that adjusting the testing inputs will not successfully attack the already trained models.